# Systematic Study on the $\alpha$-particle preformation factor in the theory of $\alpha$-decay based on the Tabular Prior-data Fitted Network (TabPFN)

### v1

@@ Abstract
A hybrid approach combining the Tabular Prior-data Fitted Network (TabPFN) with the Coulomb and Proximity Potential Model (CPPM) is developed to investigate $\alpha$-particle preformation factors $P_{\alpha}$ and their impact on $\alpha$-decay half-lives. The TabPFN model, trained on 498 nuclei, accurately learns the relationship between the properties of the nuclear structure and $P_{\alpha}$, achieving a root mean square deviation of $\sigma_{\mathrm{rms}}=0.211$. The predicted factors reveal clear odd-even staggering and shell closure effects, and exhibit a linear correlation with $Q_{\alpha}^{-1/2}$, extending the Geiger-Nuttall systematics. When incorporated into CPPM calculations, the machine learning-based $P_{\alpha}$ values significantly improve half-life predictions. The capability of the model is demonstrated through predictions for superheavy nuclei (Z=117Z=117–120), suggesting $N=184$ as a potential neutron magic number.
Usage
Secondary publications and information retrieval purposes.
Structure
[…]
@@ I Introduction
The phenomenon of $\alpha$-decay was first discovered by Rutherford in 1899. In 1928, Gamow and Condon theoretically established that $\alpha$-decay originates from the quantum tunneling effect [gamow1928quantentheorie, gurney1928wave]. Since then, $\alpha$-decay has been recognized as a pivotal tool for investigating the structure of unstable atomic nuclei, offering critical insights into their internal dynamics.
@@
Consequently, the development of a robust theoretical framework for $\alpha$-decay is of paramount importance. Advances in $\alpha$-decay theory have primarily addressed aspects including potential barrier profiles, decay energies [jin2008alpha, ni2012binding], nuclear medium effects [deng2019significant], nuclear deformation [guo2015nuclear], and the $\alpha$-particle preformation factor [zhang2009preformation, 2016Systematics]. Research on $\alpha$-decay theory has yielded numerous achievements, such as empirical formulas [viola1966nuclear, ren2012new, royer2000alpha, brown1992simple, qi2009universal], macro-microscopic models, and microscopic approaches [royer2000alpha, royer1985static, royer2010analytic, joseph2018theoretical, sun2016systematic, deng2017systematic, dong2010alpha, xu2006global].
@@
In recent years, machine learning (ML) algorithms have emerged as prominent alternative tools for investigating and predicting complex physical data, owing to their strong performance and increasingly vital role in this domain [lecun2015deep, schmidhuber2015deep]. In nuclear physics, ML techniques have been successfully applied to predict nuclear masses [carnini2020trees, wu2021nuclear, wu2020predicting], binding energies [murarka2022neutron], charge radii [wu2020calculation], $\beta$-decay lifetimes [clark2006application], and $\alpha$-decay rates [saxena2021modified]. In studies of heavy nuclei and density functional theory (DFT), ML facilitates the calibration of energy density functional (EDF) parameters [bollapragada2020optimization, goriely2014uncertainties] and enables energy predictions for over a thousand nuclides with low computational cost [lasseri2020taming]. In investigations of nuclear properties, ML is employed to assess model-related uncertainties in multi-model frameworks [connell2021does]; in nuclear reaction and neutron star studies, research focuses on the equation of state (EoS) of nuclear matter [margueron2018equationa, margueron2018equationb].
@@
Deep learning techniques have been employed in studies of $\alpha$-decay half-lives [you2024calculating, shree2025alpha, jalili2024decay, jin2023bayesian]. In Ref. [you2024calculating], an artificial neural network (ANN) was utilized to compute $\alpha$-decay half-lives; incorporating the angular momentum carried by $\alpha$ particles and the quadrupole deformation parameters of parent nuclei reduced the root-mean-square deviation (RMSD) between predictions and experimental data from 0.581 to 0.334. In Ref. [shree2025alpha], the XGBoost algorithm was adopted to predict $\alpha$-decay half-lives of superheavy nuclei, demonstrating the smallest mean square deviation compared to empirical formulas and experimental data. In Ref. [jalili2024decay], the radial basis function (RBF) kernel support vector machine (SVM) was applied to predict nuclear $\alpha$-decay half-lives; analysis of 2232 data points showed excellent performance, with root-mean-square errors (RMSE) of 0.819 and 0.352 on two datasets, comparable to other ML methods. In Ref. [jin2023bayesian], a Bayesian neural network (BNN) combined with a complex $\alpha$-decay model was proposed, leading to improved half-life predictions through modifications to the preformation factor.
@@
Within the framework of $\alpha$-decay theory, the $\alpha$-particle preformation factor characterizes the probability of $\alpha$-cluster formation on the surface of the parent nucleus. Typically, $P_{\alpha}$ is treated as an empirical constant, based on its smooth variation in open-shell regions [hodgson2003cluster]; however, this assumption may introduce inaccuracies in reproducing half-lives for nuclides near closed-shell regions. In closed-shell nuclides (e.g., those with proton number $Z=82$ and neutron number $N=126$), $P_{\alpha}$ exhibits significant variations [andreyev2013signatures, deng2015realistic], potentially leading to systematic deviations between theoretical and experimental $\alpha$-decay half-lives. Recent work by Zhang and Deng [deng2021correlation] demonstrated progress by extracting preformation factors via a semi-empirical formula, enhancing consistency between predictions and observations. The discrepancy between theoretical and experimental half-lives primarily stems from the treatment of $P_{\alpha}$, and given ML’s unique capability to address discrepancies between datasets and extract physical insights, its potential for investigating $\alpha$-particle preformation factors warrants further exploration.
@@ II.1 The extraction of experimental preformation factors $P_{\alpha}$ exp under the framework of CPPM
The $\alpha$-decay half-life of atomic nuclei is given by [qi2023systematic]:
@@
The Coulomb and Proximity Potential Model (CPPM) is employed as the unified decay theory framework, owing to its accurate description of nuclear fusion, fission, and particularly $\alpha$ decay. The assault frequency $\nu$ is set to $1.0\times 10^{22},\text{s}^{-1}$ [poenaru2011single]. The penetration probability $P$ is computed using the semiclassical Wentzel-Kramers-Brillouin (WKB) approximation as
@@
Here, $r$ is the separation between the centers of the daughter nucleus and $\alpha$ particle. The nuclear potential is described using the proximity potential formalism proposed by Blocki et al. [blocki1977proximity]:
@@
The quantity $C_{i}$ ($i=\alpha,d$) is given by $C_{i}=R_{i}\left[1-\left(\frac{b}{R_{i}}\right)^{2}\right]$, where $C_{\alpha}$ and $C_{d}$ are the matter radii of the $\alpha$ particle and daughter nucleus, respectively. The effective sharp radius $R_{i}$ ($i=\alpha,d$) is [royer2000alpha]: $R_{i}=1.28A_{i}^{1/3}-0.76+0.8A_{i}^{-1/3}$, where $A_{i}$ is the mass number of the $\alpha$ particle ($i=\alpha$) or daughter nucleus ($i=d$).
The universal function $\Phi(\xi)$ is defined as
$\phi(\xi)=\begin{cases}\frac{1}{2}(\xi-2.54)^{2}-0.0852(\xi-2.54)^{3},&\xi<1.2511,\\ -3.437\exp\left(-\frac{\xi}{0.75}\right),&\xi\geq 1.2511.\end{cases}$ (10)
with $\xi=(r-C_{\alpha}-C_{d})/b$ representing the separation between the near surfaces of the daughter nucleus and $\alpha$ particle.
@@
$V_{c}(r)=\begin{cases}\frac{Z_{\alpha}Z_{d}e^{2}}{2R}\left[3-\left(\frac{r}{R}\right)^{2}\right],&r\leq R,\\ \frac{Z_{\alpha}Z_{d}e^{2}}{r},&r>R.\end{cases}$ (11)
Here, $R=R_{d}+R_{\alpha}$ is the sum of the radii of the daughter nucleus and $\alpha$ particle, and $Z_{d}$ and $Z_{\alpha}$ are the proton numbers of the daughter nucleus and $\alpha$ particle, respectively.
The centrifugal potential $V_{l}(r)$ employs the Langer-modified form, where $l(l+1)\to\left(l+\frac{1}{2}\right)^{2}$, essential for one-dimensional problems [morehead1995asymptotics]:
$V_{l}(r)=\frac{\hbar^{2}\left(l+\frac{1}{2}\right)^{2}}{2\mu r^{2}},$ (12)
where $l$ is the orbital angular momentum carried by the emitted $\alpha$ particle. The minimum $l$ is determined by conservation of parity and angular momentum.
For the integral bounds in Eq. (6), the outer turning point radius is $R_{\text{out}}=\frac{Z_{d}Z_{\alpha}e^{2}}{2Q_{\alpha}}+\sqrt{\left(\frac{Z_{d}Z_{\alpha}e^{2}}{2Q_{\alpha}}\right)^{2}+\frac{\hbar^{2}l(l+1)}{2\mu Q_{\alpha}}}$, and the inner turning point is $R_{\text{in}}=R_{d}+R_{\alpha}$, corresponding to the separation configuration [dong2009cluster].
@@ II.2 Tabular Prior-data Fitted Network method
The Tabular Foundation Model (TabPFN) is employed—a Transformer-based architecture specifically designed for small-to-medium-sized tabular datasets (containing up to 10,000 samples and 500 features). Consider a single tabular dataset $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$ comprising $N$ training samples (table rows). Each instance $x_{i}$ is characterized by $d$ features (columns), where $d$ may vary across datasets. The label $y$ belongs to $\mathcal{C}]={1,\ldots,C}$ for classification tasks or is a numerical value for regression tasks. All attributes are assumed numerical (continuous); categorical (discrete) attributes, if present, are preprocessed using ordinal or one-hot encoding. The objective is to learn a mapping $f$ from instances to their corresponding labels. Specifically, given an unseen instance $x^{*}\in\mathbb{R}^{d}$ sampled from the same distribution as $\mathcal{D}$, the learned mapping $f$ predicts its label as $\hat{y}^{*}=f(x^{*}\mid\mathcal{D})$. A smaller discrepancy between $\hat{y}^{*}$ and the true label $y^{*}$ indicates stronger generalization capability [ye2025closer].
The complete input to TabPFN —comprising both the training set and the test instance—forms a tensor of shape $(N+1)\times(d+1)\times k$. The model alternately applies two types of self-attention: across samples (among the N+1N{+}1 instances) and across attributes (among the d+1d{+}1 dimensions), enabling in-context learning along both axes [hollmann2025accurate, ye2025closer]. Finally, the output token corresponding to the test instance’s dummy label $\tilde{y}^{*}$ is extracted and mapped either to a 10-class logit (classification) or a single-value logit (regression).
The weights of TabPFN are pretrained on diverse synthetic datasets generated via structural causal models (SCMs), with checkpoints selected according to performance on real-world datasets. Further details regarding feature preprocessing, acceleration techniques, and post-hoc ensembling are provided in Ref. [hollmann2025accurate].
In practical applications, 498 nuclear features are input to the TabPFN model, including ten parameters: mass number $A$, proton number $Z$, neutron number $N$, $\alpha$-decay energy $Q_{\alpha}$, minimum angular momentum $l$, quadrupole deformation $\beta_{2}$ of the parent nucleus [moller1993nuclear, denisov2025alpha], parity of $Z$ and NN ($Z_{p}$, $N_{p}$), and pairing effect $\delta$ [kirson2008mutual]. Leveraging the in-context learning (ICL) mechanism, the model processes the entire dataset in a single forward pass. TabPFN is pretrained on millions of synthetic datasets generated via structural causal models (SCMs). After pretraining, TabPFN can be directly applied to new real-world datasets without further training.
@@ II.4 Preformation factor and nuclear structure information
The TabPFN12 model, which exhibits optimal predictive performance, is employed to estimate the $\alpha$-particle preformation factors for 498 nuclei. The logarithmic preformation factors derived from the TabPFN12 method, along with the corresponding $\alpha$-decay energy $Q_{\alpha}$ (in MeV)—a key input to the network—are presented in Fig. 1 as functions of the proton number $Z$ and neutron number $N$. For $Z<82$ and $N<126$, $Q_{\alpha}$ increases gradually with proton and neutron numbers, whereas $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ decreases. Beyond these shell closures, both $Q_{\alpha}$ and $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ exhibit significant variations. Specifically, beyond the $Z=82$ shell closure, $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ initially rises before decreasing as $Z$ approaches the next proton magic number. A similar trend is observed for $N>126$. Evidently, the trends of $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ differ markedly across the $Z=82$ and $N=126$ closed shells, indicating that the preformation factor sensitively reflects nuclear shell effects. Thus, the preformation factor constitutes a reliable indicator for probing the stability of superheavy nuclei. Consistent with Ref. [deng2021correlation], the trend of $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ confirms that the shell effect at $^{208}\text{Pb}$ ($Z=82$, $N=126$) is stronger than that at $^{132}\text{Sn}$ ($Z=50$, $N=82$). Furthermore, the highly regular and symmetric trends exhibited by the logarithmic preformation factor and the decay energy strongly support the empirical formula proposed in Ref. [deng2021correlation].
@@
An approximately linear correlation is observed between the logarithm of the experimental $\alpha$-particle preformation factor, $\log_{10}P_{\alpha}^{\mathrm{exp}}$, and the inverse square root of the $\alpha$-decay energy, $Q_{\alpha}^{-1/2}$, in both the $N\leq 126$ and $N>126$ regions [deng2021correlation]. It is therefore pertinent to examine whether the TabPFN12 method employed in this study can adequately capture this empirical trend.Figures 2(a) and 2(b) in the upper row illustrate the relationship of $\log_{10}P_{\alpha}^{\mathrm{exp}}$ and $\log_{10}P_{\alpha}^{\mathrm{TabPFN12}}$ with $Q_{\alpha}^{-1/2}$ for nuclei with $N\leq 126$, while Figures 2(d) and 2(e) in the lower row show the corresponding results for the $N>126$ region. Each plot includes a linear regression curve along with the residual sum of squares (RSS) to quantitatively assess the goodness of fit. Furthermore, Figs. 2(c) and 2(f) display the deviation between the experimental values and the TabPFN12 predictions, i.e., $\log_{10}P_{\alpha}^{\mathrm{exp}}-\log_{10}P_{\alpha}^{\mathrm{TabPFN12}}$, as a function of $Q_{\alpha}^{-1/2}$. The results confirm a clear linear dependence between $\log_{10}P_{\alpha}^{\mathrm{exp}}$ and $Q_{\alpha}^{-1/2}$, which is corroborated by the regression parameters. Moreover, the residuals are distributed randomly around zero without any systematic trend, indicating that the TabPFN model has effectively learned the underlying physical relationship between these quantities. These findings suggest that the Geiger-Nuttall law, traditionally applied to $\alpha$-decay half-lives, may also be extended to describe the $\alpha$-particle preformation factors.

### v2

@@ Abstract
A hybrid approach combining the Tabular Prior-data Fitted Network (TabPFN) with the Coulomb and Proximity Potential Model (CPPM) is developed to investigate $\alpha$-particle preformation factors $P_{\alpha}$ and their impact on $\alpha$-decay half-lives. The TabPFN model, trained on 498 nuclei, accurately learns the relationship between nuclear structure properties and $P_{\alpha}$, achieving a root mean square deviation of $\sigma_{\mathrm{rms}}=0.211$. The predicted factors reveal clear odd-even staggering and shell closure effects, and exhibit linear correlations with both $Q_{\alpha}^{-1/2}$ and the fragmentation potential $V_{\mathrm{frag}}$. When incorporated into CPPM calculations, the machine-learning-based $P_{\alpha}$ values significantly improve half-life predictions. Similar improvements are also obtained when deformation effects are included in the potential barrier description. The capability of the model is further demonstrated through predictions for superheavy nuclei (Z=117Z=117–120), suggesting $N=184$ as a potential neutron magic number.
@@ I Introduction
The phenomenon of $\alpha$-decay was first discovered by Rutherford in 1899. In 1928, Gamow and Condon theoretically established that $\alpha$-decay originates from the quantum tunneling effect [20, 23]. Since then, $\alpha$-decay has been recognized as a pivotal tool for investigating the structure of unstable atomic nuclei, offering critical insights into their internal dynamics.
@@
Consequently, the development of a robust theoretical framework for $\alpha$-decay is of paramount importance. Advances in $\alpha$-decay theory have primarily addressed aspects including potential barrier profiles, decay energies [29, 42], nuclear medium effects [14], nuclear deformation [22], and the $\alpha$-particle preformation factor [65, 12]. Furthermore, nuclear deformation may influence the barrier penetration probability in $\alpha$ decay, since the quadrupole deformation of the Coulomb barrier can modify both the height and the width of the potential barrier [10]. Research on $\alpha$-decay theory has yielded numerous achievements, such as empirical formulas [55, 47, 49, 5, 45], macro-microscopic models, and microscopic approaches [49, 48, 50, 30, 54, 16, 18, 60].
@@
In recent years, machine learning (ML) algorithms have emerged as prominent alternative tools for investigating and predicting complex physical data, owing to their strong performance and increasingly vital role in this domain [35, 52]. In nuclear physics, ML techniques have been successfully applied to predict nuclear masses [6, 58, 59], binding energies [41], charge radii [57], $\beta$-decay lifetimes [7], and $\alpha$-decay rates [51]. In studies of heavy nuclei and density functional theory (DFT), ML facilitates the calibration of energy density functional (EDF) parameters [4, 21] and enables energy predictions for over a thousand nuclides with low computational cost [34]. In investigations of nuclear properties, ML is employed to assess model-related uncertainties in multi-model frameworks [8]; in nuclear reaction and neutron star studies, research focuses on the equation of state (EoS) of nuclear matter [37, 38].
@@
Deep learning techniques have been employed in studies of $\alpha$-decay half-lives [64, 53, 27, 28]. In Ref. [64], an artificial neural network (ANN) was utilized to compute $\alpha$-decay half-lives; incorporating the angular momentum carried by $\alpha$ particles and the quadrupole deformation parameters of parent nuclei reduced the root-mean-square deviation (RMSD) between predictions and experimental data from 0.581 to 0.334. In Ref. [53], the XGBoost algorithm was adopted to predict $\alpha$-decay half-lives of superheavy nuclei, demonstrating the smallest mean square deviation compared to empirical formulas and experimental data. In Ref. [27], the radial basis function (RBF) kernel support vector machine (SVM) was applied to predict nuclear $\alpha$-decay half-lives; analysis of 2232 data points showed excellent performance, with root-mean-square errors (RMSE) of 0.819 and 0.352 on two datasets, comparable to other ML methods. In Ref. [28], a Bayesian neural network (BNN) combined with a complex $\alpha$-decay model was proposed, leading to improved half-life predictions through modifications to the preformation factor.
@@
Within the framework of $\alpha$-decay theory, the $\alpha$-particle preformation factor characterizes the probability of $\alpha$-cluster formation on the surface of the parent nucleus. Typically, $P_{\alpha}$ is treated as an empirical constant, based on its smooth variation in open-shell regions [25]; however, this assumption may introduce inaccuracies in reproducing half-lives for nuclides near closed-shell regions. In closed-shell nuclides (e.g., those with proton number $Z=82$ and neutron number $N=126$), $P_{\alpha}$ exhibits significant variations [1, 13], potentially leading to systematic deviations between theoretical and experimental $\alpha$-decay half-lives. Recent work by Zhang and Deng [15] demonstrated progress by extracting preformation factors via a semi-empirical formula, enhancing consistency between predictions and observations. The discrepancy between theoretical and experimental half-lives primarily stems from the treatment of $P_{\alpha}$, and given ML’s unique capability to address discrepancies between datasets and extract physical insights, its potential for investigating $\alpha$-particle preformation factors warrants further exploration.
@@ II.1 The extraction of experimental preformation factors $P_{\alpha}$ exp under the framework of CPPM
The $\alpha$-decay half-life of atomic nuclei is given by [46]:
@@
The Coulomb and Proximity Potential Model (CPPM) is employed as the unified decay theory framework, owing to its accurate description of nuclear fusion, fission, and particularly $\alpha$ decay. The assault frequency $\nu$ is set to $1.0\times 10^{22},\text{s}^{-1}$ [43]. The penetration probability $P$ is computed using the semiclassical Wentzel-Kramers-Brillouin (WKB) approximation as
@@
Here, $r$ is the separation between the centers of the daughter nucleus and $\alpha$ particle. The nuclear potential is described using the proximity potential formalism proposed by Blocki et al. [3]:
@@
The quantity $C_{i}$ ($i=\alpha,d$) is given by $C_{i}=R_{i}\left[1-\left(\frac{b}{R_{i}}\right)^{2}\right]$, where $C_{\alpha}$ and $C_{d}$ are the matter radii of the $\alpha$ particle and daughter nucleus, respectively. The effective sharp radius $R_{i}$ ($i=\alpha,d$) is [49]: $R_{i}=1.28A_{i}^{1/3}-0.76+0.8A_{i}^{-1/3}$, where $A_{i}$ is the mass number of the $\alpha$ particle ($i=\alpha$) or daughter nucleus ($i=d$).
The universal function $\Phi(\xi)$ is defined as {align} ϕ(ξ) = {12(ξ- 2.54)^2 - 0.0852(ξ- 2.54)^3, ξ¡ 1.2511,
-3.437 exp(-ξ0.75), ξ≥1.2511. with $\xi=(r-C_{\alpha}-C_{d})/b$ representing the separation between the near surfaces of the daughter nucleus and $\alpha$ particle.
@@
Here,
[…]
@@ II.2 Tabular Prior-data Fitted Network method
The Tabular Foundation Model (TabPFN) is employed—a Transformer-based architecture specifically designed for small-to-medium-sized tabular datasets (containing up to 10,000 samples and 500 features). Consider a single tabular dataset $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$ comprising $N$ training samples (table rows). Each instance $x_{i}$ is characterized by $d$ features (columns), where $d$ may vary across datasets. The label $y$ belongs to $\mathcal{C}]={1,\ldots,C}$ for classification tasks or is a numerical value for regression tasks. All attributes are assumed numerical (continuous); categorical (discrete) attributes, if present, are preprocessed using ordinal or one-hot encoding. The objective is to learn a mapping $f$ from instances to their corresponding labels. Specifically, given an unseen instance $x^{*}\in\mathbb{R}^{d}$ sampled from the same distribution as $\mathcal{D}$, the learned mapping $f$ predicts its label as $\hat{y}^{*}=f(x^{*}\mid\mathcal{D})$. A smaller discrepancy between $\hat{y}^{*}$ and the true label $y^{*}$ indicates stronger generalization capability [63].
The complete input to TabPFN —comprising both the training set and the test instance—forms a tensor of shape $(N+1)\times(d+1)\times k$. The model alternately applies two types of self-attention: across samples (among the N+1N{+}1 instances) and across attributes (among the d+1d{+}1 dimensions), enabling in-context learning along both axes [26, 63]. Finally, the output token corresponding to the test instance’s dummy label $\tilde{y}^{*}$ is extracted and mapped either to a 10-class logit (classification) or a single-value logit (regression).
The weights of TabPFN are pretrained on diverse synthetic datasets generated via structural causal models (SCMs), with checkpoints selected according to performance on real-world datasets. Further details regarding feature preprocessing, acceleration techniques, and post-hoc ensembling are provided in Ref. [26].
In practical applications, 498 nuclear features are input to the TabPFN model, including ten parameters: mass number $A$, proton number $Z$, neutron number $N$, $\alpha$-decay energy $Q_{\alpha}$, minimum angular momentum $l$, quadrupole deformation $\beta_{2}$ of the parent nucleus [39, 17], parity of $Z$ and NN ($Z_{p}$, $N_{p}$), and pairing effect $\delta$ [31]. Leveraging the in-context learning (ICL) mechanism, the model processes the entire dataset in a single forward pass. TabPFN is pretrained on millions of synthetic datasets generated via structural causal models (SCMs). After pretraining, TabPFN can be directly applied to new real-world datasets without further training.
@@ II.3 Effect of Ten Nuclear Physics Descriptors on Alpha Decay Preformation Factor with TabPFN Model
To further examine the role of nuclear deformation, the $\alpha$-particle preformation factors were also extracted using a deformed potential that combines the Coulomb interaction and the Woods–Saxon nuclear potential following the approach of Ref. [9]. The resulting preformation factors were subsequently predicted using the same TabPFN12 model. In this case, the obtained root-mean-square deviation is $\sigma_{\mathrm{RMS}}=0.232$, which remains close to the value obtained with the CPPM-based extraction. This agreement indicates that the TabPFN framework provides stable predictions for the $\alpha$-preformation factors even when different theoretical potentials are adopted.
@@ II.4 Preformation factor and nuclear structure information
The TabPFN12 model, which exhibits optimal predictive performance, is employed to estimate the $\alpha$-particle preformation factors for 498 nuclei. The logarithmic preformation factors derived from the TabPFN12 method, along with the corresponding $\alpha$-decay energy $Q_{\alpha}$ (in MeV)—a key input to the network—are presented in Fig. 1 as functions of the proton number $Z$ and neutron number $N$. For $Z<82$ and $N<126$, $Q_{\alpha}$ increases gradually with proton and neutron numbers, whereas $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ decreases. Beyond these shell closures, both $Q_{\alpha}$ and $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ exhibit significant variations. Specifically, beyond the $Z=82$ shell closure, $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ initially rises before decreasing as $Z$ approaches the next proton magic number. A similar trend is observed for $N>126$. Evidently, the trends of $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ differ markedly across the $Z=82$ and $N=126$ closed shells, indicating that the preformation factor sensitively reflects nuclear shell effects. Thus, the preformation factor constitutes a reliable indicator for probing the stability of superheavy nuclei. Consistent with Ref. [15], the trend of $\log_{10}P_{\alpha}^{\text{TabPFN12}}$ confirms that the shell effect at $^{208}\text{Pb}$ ($Z=82$, $N=126$) is stronger than that at $^{132}\text{Sn}$ ($Z=50$, $N=82$). Furthermore, the highly regular and symmetric trends exhibited by the logarithmic preformation factor and the decay energy strongly support the empirical formula proposed in Ref. [15].
@@
An approximately linear correlation is observed between the logarithm of the experimental $\alpha$-particle preformation factor, $\log_{10}P_{\alpha}^{\mathrm{exp}}$, and the inverse square root of the $\alpha$-decay energy, $Q_{\alpha}^{-1/2}$, in both the $N\leq 126$ and $N>126$ regions [15]. It is therefore pertinent to examine whether the TabPFN12 method employed in this study can adequately capture this empirical trend. Figures Figure 2 (a) and  2 (b) in the upper row illustrate the relationship of $\log_{10}P_{\alpha}^{\mathrm{exp}}$ and $\log_{10}P_{\alpha}^{\mathrm{TabPFN12}}$ with $Q_{\alpha}^{-1/2}$ for nuclei with $N\leq 126$, while Figures 2 (d) and 2 (e) in the lower row show the corresponding results for the $N>126$ region. Each plot includes a linear regression curve along with the residual sum of squares (RSS) to quantitatively assess the goodness of fit. Furthermore, Figs. 2 (c) and 2 (f) display the deviation between the experimental values and the TabPFN12 predictions, i.e., $\log_{10}P_{\alpha}^{\mathrm{exp}}-\log_{10}P_{\alpha}^{\mathrm{TabPFN12}}$, as a function of $Q_{\alpha}^{-1/2}$. The results confirm a clear linear dependence between $\log_{10}P_{\alpha}^{\mathrm{exp}}$ and $Q_{\alpha}^{-1/2}$, which is corroborated by the regression parameters. Moreover, the residuals are distributed randomly around zero without any systematic trend, indicating that the TabPFN model has effectively learned the underlying physical relationship between these quantities. These findings suggest that the Geiger-Nuttall law, traditionally applied to $\alpha$-decay half-lives, may also be extended to describe the $\alpha$-particle preformation factors.
Besides the correlation with $Q_{\alpha}$, the $\alpha$-particle preformation probability is also expected to depend on the fragmentation potential $V_{\mathrm{frag}}$, which characterizes the energy cost of forming an $\alpha$ cluster in the parent nucleus [44, 11]. To explore this dependence, the relationship between $\log_{10}P_{\alpha}$ and $V_{\mathrm{frag}}$ is examined for the same set of 498 nuclei. Figure 3(a) shows the correlation between the experimentally extracted preformation factors and the fragmentation potential. A clear negative linear trend is observed, indicating that larger fragmentation potentials correspond to smaller $\alpha$-cluster formation probabilities. The linear regression gives
$\log_{10}P_{\alpha}^{\mathrm{exp}}=-0.119V_{\mathrm{frag}}-0.119.$ (30)
For comparison, Fig. 3(b) presents the corresponding relationship for the TabPFN12-predicted preformation factors. A very similar linear dependence is obtained,
$\log_{10}P_{\alpha}^{\mathrm{TabPFN12}}=-0.112V_{\mathrm{frag}}-0.221.$ (31)
The close agreement between the slopes of the experimental and TabPFN12-derived correlations demonstrates that the machine learning model successfully captures the underlying physical dependence of the $\alpha$-cluster preformation probability on the fragmentation potential. Physically, a larger $V_{\mathrm{frag}}$ implies a higher energy cost for cluster formation, leading to a reduced probability for $\alpha$ preformation. This result further confirms that the TabPFN model not only reproduces experimental trends but also retains meaningful nuclear-structure information.
Figure 3: Correlation between the fragmentation potential $V_{\mathrm{frag}}$ and the logarithm of the $\alpha$-particle preformation factor for the 498 nuclei. (a) Experimentally extracted preformation factors $\log_{10}P_{\alpha}^{\mathrm{exp}}$. (b) Preformation factors predicted by the TabPFN12 model $\log_{10}P_{\alpha}^{\mathrm{TabPFN12}}$. The red lines represent linear fits to the data.